likeabossssssss:

A teacher randomly selects one student from a group of students where 12 are wearing a black shirt, 8 are wearing a red shirt, and 12 are wearing a blue shirt. What is the probability that the selected student is wearing a blue shirt? Enter your answer as a fraction, in simplified form, in the box.

likeabossssssss:

@KjSaif @SkyVoltage43

Justin234:

Probability=number of specific outcomes/number of possible outcomes... P(B)=B/T=12/(12+12+8) P(B)=12/32 P(B)=3/8

Justin234:

Like a boss *Dab*

Justin234:

Hope this helps

SkyVoltage43:

and like a cheater *dab*

KjSaif:

^^

SkyVoltage43:

no medals for u fine sir

Justin234:

eh no its called Answering

KjSaif:

more like copying and pasting for u man.

SkyVoltage43:

its called copying and pasting answers from brainly

Justin234:

But isnt what other ppl gunna do here after were done with this question :/

KjSaif:

U stole that from *COUGH* brainly *COUGH*

Justin234:

and its right tho....

KjSaif:

really? are you sure? Doubt by the rating man.

SkyVoltage43:

@Ultrilliam

likeabossssssss:

what is going on?

Justin234:

idk My answer was rite so idk what these freaks are doing

KjSaif:

he stole the answer from brainly

Justin234:

Not stole used

likeabossssssss:

yea

Justin234:

Der is nothing wrong with useing brainly

likeabossssssss:

ok what ever stop spamin my post k

SkyVoltage43:

mkay

Justin234:

So no Medal?

SkyVoltage43:

nope